# BrainECHO: Semantic Brain Signal Decoding through Vector-Quantized Spectrogram Reconstruction for Whisper-Enhanced Text Generation

### v1

@@ Abstract
Recent advances in decoding language from brain signals (EEG and MEG) have been significantly driven by pre-trained language models, leading to remarkable progress on publicly available non-invasive EEG/MEG datasets. However, previous works predominantly utilize teacher forcing during text generation, leading to significant performance drops without its use. A fundamental issue is the inability to establish a unified feature space correlating textual data with the corresponding evoked brain signals. Although some recent studies attempt to mitigate this gap using an audio-text pre-trained model, Whisper, which is favored for its signal input modality, they still largely overlook the inherent differences between audio signals and brain signals in directly applying Whisper to decode brain signals. To address these limitations, we propose a new multi-stage strategy for semantic brain signal decoding via vEctor-quantized speCtrogram reconstruction for WHisper-enhanced text generatiOn, termed BrainECHO. Specifically, BrainECHO successively conducts: 1) Discrete autoencoding of the audio spectrogram; 2) Brain-audio latent space alignment; and 3) Semantic text generation via Whisper finetuning. Through this autoencoding–alignment–finetuning process, BrainECHO outperforms state-of-the-art methods under the same data split settings on two widely accepted resources: the EEG dataset (Brennan) and the MEG dataset (GWilliams). The innovation of BrainECHO, coupled with its robustness and superiority at the sentence, session, and subject-independent levels across public datasets, underscores its significance for language-based brain-computer interfaces.
Introduction
@@
Decoding text from brain activity, such as electroencephalography (EEG) and magnetoencephalography (MEG), is a critical and frontier research topic, that can provide a foundation for language-based brain-computer interfaces (BCI) by enabling direct text input through brain signals. In the long term, accurate real-time translation of human brain signals can promote the widespread application of BCI technology in medicine, assistive technology, and entertainment, bringing new possibilities to human life.
@@
With the rapid developments in natural language processing (NLP), automatic speech recognition (ASR), and other fields, researchers have leveraged the powerful language understanding and generating capabilities of pretrained large language models (LLMs) for neural decoding tasks (Wang and Ji 2022; Duan et al. 2024; Yang et al. 2024a, b), making it possible to accurately decode text stimuli from non-invasive signals. EEG-to-Text (Wang and Ji 2022) is the first work to decode open-vocabulary tokens from encoded word-level EEG rhythm features with the pretrained large model BART (Lewis et al. 2020). Furthermore, DeWave (Duan et al. 2024) used sentence-level raw EEG signals to perform EEG-to-text decoding without eye movement event markers.
@@
Later on, several BART-based methods (Xi et al. 2023; Feng et al. 2023; Amrani, Micucci, and Napoletano 2024) were introduced, predominantly employing a pretraining-finetuning paradigm. These methods first align EEG representations with pretrained text embeddings before feeding them into BART for finetuning. Although these approaches have yielded impressive results, they rely on a teacher-forcing generation strategy, wherein the model depends on the ground truth preceding text during each token prediction. This setting does not accurately reflect the model’s performance in real-world scenarios. These methods show poor decoding performance without teacher forcing.
@@
To address this limitation, NeuSpeech (Yang et al. 2024a) and MAD (Yang et al. 2024b) treat raw MEG signals as a specialized form of speech, transforming MEG signals and feeding them into a pre-trained Whisper model (Radford et al. 2023), which is trained on large-scale audio-text pairs, for end-to-end text decoding without teacher forcing. However, these approaches primarily focus on mapping continuous brain signals to discrete text without compressing the signals into discrete representations, thereby limiting the model’s decoding accuracy and generalization capabilities.
Extensive researches in speech recognition (Zhang et al. 2023a; Puvvada et al. 2024) demonstrate that discrete representations preserve more semantic information for translation compared to conventional speech features like Fbank, thanks to their carefully designed self-supervised learning paradigms. While DeWave (Duan et al. 2024) aligns discrete representations of input EEG signals and text, it assumes a chronological order for the discrete token sequence, requiring a highly capable feature extractor. Considering the natural temporal alignment between audio-evoked brain signals and audio stimuli, aligning raw signals and speech within a discrete space leverages implicit temporal properties, thereby reducing the difficulty of training.
Therefore, we propose a novel multi-stage semantic decoding framework for EEG/MEG brain signals, aurally evoked by semantic audio, through vEctor-quantized speCtrogram reconstruction for WHisper-enhanced text generatiOn, termed BrainECHO. Specifically, BrainECHO executes the following steps: 1) Discrete autoencoding of the audio spectrogram, particularly employing codebook-based vector quantization, to establish a pre-warmed representation space that facilitates Mel spectrogram reconstruction; 2) Brain-audio latent space alignment, utilizing a brain encoder and pre-warmed quantizer and decoder to reconstruct the evoked brain signal’s Mel spectrogram; 3) Semantic text generation, achieved through AdaLoRA-based finetuning of the pre-trained Whisper model, with the reconstructed Mel spectrogram as input. The overall three-stage (autoencoding, alignment, finetuning) training process of the proposed BrainECHO is illustrated in Figure 1. We validate the performance of BrainECHO using two different public audio-evoked brain signal datasets: Brennan, which contains EEG data, and GWilliams, which contains MEG data. The principal contributions of our work are summarized below:
  • The proposed BrainECHO framework overcomes the current flaw in EEG/MEG-to-text approaches that mistakenly rely on the teacher-forcing strategy. It achieves semantic decoding with significantly improved results compared to using Gaussian noise as the input.
  • We propose breaking down the EEG/MEG-to-text task into a multi-stage strategy to mitigate the biases induced by the overwhelming capabilities of large language models, while still leveraging their pre-trained knowledge, specifically utilizing Whisper in our work.
  • We introduce vector-quantized discrete representations to enhance the model’s efficiency, achieving state-of-the-art (SOTA) performance on EEG and MEG datasets. Specifically, we evaluate BrainECHO across various data split scenarios, which are neglected in prior research.
@@
Figure 1: Learning process overview of our proposed BrainECHO framework. BrainECHO follows a three-stage autoencoding–alignment–finetuning paradigm: Autoencoding stage is used to warm up the Mel spectrogram reconstruction by employing a codebook-based quantizer to enhance generalizability and robustness. This stage especially focuses on exploiting discrete representations. Alignment stage reconstructs the Mel spectrogram from the corresponding aurally evoked brain signals. This involves designing a new brain encoder that integrates with the warmed-up quantizer and decoder from the first stage. Finetuning stage leverages the capabilities of the pre-trained Whisper model to achieve audio-text translation.
Figure 2: (a) Overview of the BrainECHO model framework. BrainECHO utilizes a three-stage training paradigm consisting of Mel spectrogram autoencoding, brain-audio latent space alignment and Whisper finetuning. $C$, $T_{\varepsilon}$ denotes numbers of raw wave channels and timestamps, respectively. (b) Details of the Brain Encoder, which converts raw EEG/MEG signals into latent representations. $d$ represents the dimension of hidden states and TS Conv stands for Spatio-Temporal Convolution Networks. More details of Conformer are provided in the Appendix.
Related Works
Non-invasive brain signals such as EEG and MEG offer significant advantages over invasive alternatives, particularly in terms of safety and cost-effectiveness. Significant progress has been made in decoding text from non-invasive signals.
Closed-Vocabulary Neural Decoding
@@
Ghazaryan et al. (Ghazaryan et al. 2023) utilized Word2vec to decode 60 nouns from MEG recordings. Meta (Défossez et al. 2023) developed a model employing wav2vec 2.0 (Baevski et al. 2020) and contrastive learning to decode speech from 3-second EEG/MEG signals. However, these methods are constrained to closed-vocabulary tasks, restricting their applicability in open-vocabulary text generation.
Decoder-Only Architectures for Open-Vocabulary Brain-to-Text Decoding
@@
Recent advancements have leveraged the powerful understanding and generation capabilities of pretrained models, particularly LLMs, to extend vocabulary from closed to open. In decoder-only architectures, some researchers have aligned brain signals with text to guide pretrained generative models like GPT in text generation. For example, Zhao et al.(Zhao et al. 2024) mapped fMRI data to text embeddings to iteratively guide GPT-2 in generating text. Similarly, Chen et al.(Chen et al. 2024) used text-aligned fMRI representations as prompts for GPT-2 to decode language information.
Seq2seq Architectures for Open-Vocabulary Brain-to-Text Decoding
@@
Wang et al.(Wang and Ji 2022) fed transformed word-level EEG rhythm features into a pretrained BART model to decode open-vocabulary tokens. Duan et al.(Duan et al. 2024) integrated discrete EEG encodings with text-EEG contrastive alignment to mitigate individual variability in brain activity. However, these BART-based methods rely on teacher forcing during inference. Furthermore, as Jo et al. (Jo et al. 2024) demonstrated, their performance on noisy data is comparable to that on EEG data, suggesting that these models may simply memorize the training data. Recently, NeuSpeech (Yang et al. 2024a) directly fed raw MEG signals into a modified, pretrained Whisper model for text decoding without teacher forcing. Furthermore, MAD (Yang et al. 2024b) introduced MEG-speech alignment loss to decode sentences not present in the training data. However, these Whisper-based methods do not utilize discrete representations of the original signals to enhance the model’s generalization capabilities. Our work integrates brain-audio discretization and alignment, aiming to predict high-quality Mel spectrograms from brain signals that align with Whisper’s input format. Leveraging Whisper’s advanced speech recognition abilities, our approach generates sentences that closely mirror the original text.
Method
Task Definition
@@
Given the raw EEG/MEG $E$, text content $T$, and corresponding audio stimuli $A$, the experimental data can be divided into a series of sentence-level EEG/MEG-text-speech pairs $\left\langle\varepsilon,t,a\right\rangle$ according to $T$. Our goal is to decode the corresponding open-vocabulary tokens $t$ from the brain signal $\varepsilon$, with $a$ serving as auxiliary information.
Model Architecture
@@
Unlike the multi-task joint training employed in MAD (Yang et al. 2024b), BrainECHO adopts a three-stage training process. This method reduces resource consumption at each training step and facilitates the prediction of high-quality, high-resolution Mel spectrograms from brain signals. Specifically, we extend the spectrogram duration from 3 seconds, as used in  (Défossez et al. 2023; Yang et al. 2024b), to over 10 seconds, enabling sentence-level rather than segment-level brain-to-text translation, thereby preserving the semantics of the original sentences. The details of the model are shown in Figure 2. The following sections will detail each training stage.
Discrete Autoencoding of Audio Spectrogram
@@
Van den Oord et al. introduced the Vector Quantized-Variational AutoEncoder (VQ-VAE) (Van Den Oord, Vinyals et al. 2017) to learn discrete latent representations of audio, video, and other data types. Building on this approach, several studies (Li, Jia, and Chiu 2023; Sadok, Leglaive, and Séguier 2023; Yang et al. 2023) have explored representing Mel spectrograms using discrete tokens to capture phoneme-like information. Inspired by these methods, our first stage involves autoencoding Mel spectrograms, with the purpose of obtaining a discrete representation space that is conducive to Mel reconstruction. Specifically, given a spectrogram $m\in\mathbb{R}^{T_{m}\times F_{m}}$, the audio encoder $Enc$ first converts it into a feature map $z_{m}=Enc(m)\in\mathbb{R}^{t_{m}\times f_{m}\times D}$ , where $T_{m}$, $F_{m}$ and $D$ denote the number of time frames, frequency bins and latent channels, respectively. The spectrogram is generated by the Whisper Processor, enabling text decoding from the reconstructed spectrogram using Whisper’s encoder-decoder architecture. Then, $z_{m}$ is processed by a vector quantizer $Q$. Specifically, each latent embedding $z_{m}^{ij}\in\mathbb{R}^{D}$ ($1\leq i\leq t_{m},1\leq j\leq f_{m}$) is replaced by the nearset vector $z_{q}^{ij}$ from a codebook $\mathbb{C}\in\mathbb{R}^{N\times D}$, which consists of $N$ learnable D-dimensional vectors. Formally, this process is expressed as follows:
@@
$Q(z_{m}^{ij})=z_{q}^{ij}=c_{k},\text{ where }k=\mathop{\arg\min}\limits_{k\in\{1,2,...,N\}}\left\|z_{m}^{ij}-c_{k}\right\|_{2}.$ (1)
@@
The reconstructed spectrogram is then obtained by the audio decoder $Dec$ as: $\hat{m}=Dec(z_{q})$. The encoder and decoder are both composed of ResUNet blocks (Kong et al. 2021). The training objective at this stage is defined as follows:
@@
Brain-Audio Latent Space Alignment
In the second stage, we freeze all the modules pre-trained in the previous stage and train a brain encoder to convert raw EEG/MEG signals $\varepsilon$ into latent features $z_{\varepsilon}$. The brain encoder utilizes a Conformer-based architecture (Song et al. 2022), which begins with Spatio-Temporal Convolutional Networks to process the input signals into a one-dimensional embedding sequence. The spatial convolutional layer reduces the number of input signal channels to one, while the temporal convolutional layers downsample the time dimension. This sequence is then added to learnable position embeddings and fed into a stack of Transformer encoder blocks. Linear layers and 2D convolutional networks subsequently transform the EEG/MEG features into representations matching the shape of $z_{m}$. Similarly, $z_{\varepsilon}$ is input into the frozen quantizer $Q$ and audio decoder $Dec$ to predict the corresponding Mel spectrogram $m$. Additionally, we align the representations of the Mel spectrogram and raw signals in the latent space. Notably, we employ a unified codebook to leverage pre-warmed discrete acoustic tokens for representing brain activity. The introduction of vector quantization enhances the stability and generalization of the Mel spectrogram reconstruction from brain signals, thereby improving the performance of subsequent text decoding. Formally, the loss for stage 2 is designed as follows:

### v2

@@ Abstract
Current EEG/MEG-to-text decoding systems suffer from three key limitations: (1) reliance on teacher-forcing methods, which compromises robustness during inference, (2) sensitivity to session-specific noise, hindering generalization across subjects, and (3) misalignment between brain signals and linguistic representations due to pre-trained language model over-dominance. To overcome these challenges, we propose BrainECHO (Brain signal decoding via vEctor-quantized speCtrogram reconstruction for WHisper-enhanced text generatiOn), a multi-stage framework that employs decoupled representation learning to achieve state-of-the-art performance on both EEG and MEG datasets. Specifically, BrainECHO consists of three stages: (1) Discrete autoencoding, which transforms continuous Mel spectrograms into a finite set of high-quality discrete representations for subsequent stages. (2) Frozen alignment, where brain signal embeddings are mapped to corresponding Mel spectrogram embeddings in a frozen latent space, effectively filtering session-specific noise through vector-quantized reconstruction, yielding a 3.65% improvement in BLEU-4 score. (3) Constrained decoding fine-tuning, which leverages the pre-trained Whisper model for audio-to-text translation, balancing signal adaptation with knowledge preservation, and achieving 74%-89% decoding BLEU scores without excessive reliance on teacher forcing. BrainECHO demonstrates robustness across sentence, session, and subject-independent conditions, passing Gaussian noise tests and showcasing its potential for enhancing language-based brain-computer interfaces.
BrainECHO: Semantic Brain Signal Decoding through Vector-Quantized Spectrogram Reconstruction for Whisper-Enhanced Text Generation
Jilong Li1, Zhenxi Song1thanks: Corresponding author, Jiaqi Wang1,2, Meishan Zhang1, Honghai Liu1, Min Zhang1, Zhiguo Zhang1thanks: Corresponding author 1Harbin Institute of Technology, Shenzhen, China 2Peng Cheng Laboratory, China Correspondence: {songzhenxi, zhiguozhang}@hit.edu.cn
1 Introduction
@@
Decoding text from brain activity, such as electroencephalography (EEG) and magnetoencephalography (MEG), is a critical and frontier research topic that can provide a foundation for language-based brain-computer interfaces (BCI) by enabling direct text input through brain signals. In the long term, accurate real-time translation of human brain signals can promote the widespread application of BCI technology in medicine, assistive technology, and entertainment, bringing new possibilities to human life.
@@
With the rapid developments in natural language processing (NLP), automatic speech recognition (ASR), and other fields, researchers have leveraged the powerful language understanding and generating capabilities of pretrained large language models (LLMs) for neural decoding tasks (Wang and Ji, 2022; Duan et al., 2024; Yang et al., 2024b, c), making it possible to accurately decode text stimuli from non-invasive signals. EEG-to-Text (Wang and Ji, 2022) is the first work to decode open-vocabulary tokens from encoded word-level EEG rhythm features with the pretrained large model BART (Lewis et al., 2020). Furthermore, DeWave (Duan et al., 2024) used sentence-level raw EEG signals to perform EEG-to-text decoding without eye movement event markers.
@@
Later on, several BART-based methods (Xi et al., 2023; Feng et al., 2023; Amrani et al., 2024) were introduced, predominantly employing a pretraining-finetuning paradigm. These methods first align EEG representations with pretrained text embeddings before feeding them into BART for finetuning. Although these approaches have yielded impressive results, they rely on a teacher-forcing generation strategy, wherein the model depends on the ground truth preceding text during each token prediction. This setting does not accurately reflect the model’s performance in real-world scenarios. These methods show poor decoding performance without teacher forcing.
@@
To address this limitation, NeuSpeech (Yang et al., 2024b) and MAD (Yang et al., 2024c) treat raw MEG signals as a specialized form of speech, transforming MEG signals and feeding them into a pre-trained Whisper model (Radford et al., 2023), which is trained on large-scale audio-text pairs, for end-to-end text decoding without teacher forcing. However, these approaches primarily focus on mapping continuous brain signals to discrete text without compressing the signals into discrete representations, thereby limiting the model’s decoding accuracy and generalization capabilities.
@@
Figure 1: Overview of the BrainECHO framework learning process, illustrated through a simplified conceptual diagram for enhanced understanding. BrainECHO follows a three-stage autoencoding–alignment–finetuning paradigm to achieve decoupled representation learning: Autoencoding stage is used to warm up the Mel spectrogram reconstruction by employing a codebook-based quantizer to enhance generalizability and robustness. This stage especially focuses on exploiting discrete representations. Alignment stage reconstructs the Mel spectrogram from the corresponding aurally evoked brain signals. This involves designing a new brain encoder that integrates with the warmed-up quantizer and decoder from the first stage. Finetuning stage leverages the capabilities of the pre-trained Whisper model to achieve audio-text translation.
In brain-to-text decoding, introducing discrete representation helps solve two fundamental challenges. First, the codebook’s discrete latent space serves as a modality-invariant interface, enabling seamless alignment between brain signals and text tokens. This avoids the "distribution shift" problem in end-to-end continuous-to-discrete mapping, which can lead to spurious correlations (Shirakawa et al., 2024). Second, EEG/MEG signals are inherently contaminated by physiological artifacts (e.g., muscle movements, ocular noise) and session-specific variability (e.g., electrode impedance shifts). By discretizing brain signals into a semantic-pruned codebook, vector quantization acts as a sparsity-inducing filter that discards task-irrelevant embeddings. This mechanism is analogous to noise suppression in VQ-VAE-based models (Razavi et al., 2019).
Therefore, we propose a novel multi-stage semantic decoding framework for EEG/MEG brain signals, aurally evoked by semantic audio, through vEctor-quantized speCtrogram reconstruction for WHisper-enhanced text generatiOn, termed BrainECHO. The overall three-stage (autoencoding, alignment, finetuning) training process of the proposed BrainECHO is illustrated in Figure 1. We validate the performance of BrainECHO using two different public audio-evoked brain signal datasets: Brennan, which contains EEG data, and GWilliams, which contains MEG data. The principal contributions of our work are summarized below:
  • The proposed BrainECHO framework addresses EEG/MEG-to-text limitations of teacher-forcing dependency and poor Gaussian noise generalization (Jo et al., 2024; Wang and Ji, 2022), achieving SOTA performance on EEG and MEG benchmarks (Brennan and Hale, 2019; Gwilliams et al., 2023). Its robustness is further validated through novel subject/session-independent data splits, addressing a critical gap in prior research.
  • Unlike recent non-teacher-forcing methods (Yang et al., 2024b, c) that directly fine-tune LLMs, BrainECHO mitigates LLM overfitting risks through a multi-stage training strategy, effectively balancing noise suppression in brain signals with preservation of pre-trained linguistic knowledge.
  • By introducing a quantized codebook for discrete brain-signal representation—contrary to continuous latent spaces in prior work—BrainECHO filters session-specific noise and captures subject-invariant semantics, achieving SOTA cross-subject generalization.
2 Related Works
@@
Non-invasive brain signals such as EEG and MEG offer significant advantages over invasive alternatives, particularly in terms of safety and cost-effectiveness. Considerable progress has been made in decoding text from noninvasive signals. Ghazaryan et al. (Ghazaryan et al., 2023) utilized Word2vec to decode 60 nouns from MEG recordings. Meta (Défossez et al., 2023) developed a model that uses wav2vec 2.0 (Baevski et al., 2020) and contrastive learning to decode speech from 3-second EEG/MEG signals. However, these methods are restricted to decoding a small set of words or segments, restricting their applicability to open-vocabulary text generation.
2.1 Decoder-Only Models for Brain-to-Text
@@
Recent advancements have leveraged the powerful understanding and generation capabilities of pretrained models, particularly LLMs, to extend vocabulary from closed to open. In decoder-only architectures, some researchers have aligned brain signals with text to guide pretrained generative models in text generation. For example, Tang et al. (Tang et al., 2023) and Zhao et al. (Zhao et al., 2024) mapped fMRI data to text embeddings to iteratively guide GPT-2 in generating text. Similarly, Chen et al. (Chen et al., 2024a) used text-aligned fMRI representations as prompts for GPT-2 to decode language information.
2.2 Seq2Seq Models for Brain-to-Text
@@
Wang et al.(Wang and Ji, 2022) fed transformed word-level EEG rhythm feature into a pretrained BART model to decode open-vocabulary tokens. Duan et al.(Duan et al., 2024) integrated discrete EEG encodings with text-EEG contrastive alignment to mitigate individual variability in brain activity. However, these BART-based methods rely on teacher forcing during inference. Furthermore, as Jo et al. (Jo et al., 2024) demonstrated, their performance on noisy data is comparable to that on EEG data, suggesting that these models may simply memorize the training data. Recently, NeuSpeech (Yang et al., 2024b) directly fed raw MEG signals into a modified, pretrained Whisper model for text decoding without teacher forcing. Furthermore, MAD (Yang et al., 2024c) introduced MEG-speech alignment loss to decode sentences not present in the training data. However, these Whisper-based methods do not utilize discrete representations of the original signals to enhance the model’s generalization capabilities. Our work integrates brain-audio discretization and alignment, aiming to predict high-quality Mel spectrograms from brain signals that align with Whisper’s input format. Leveraging Whisper’s advanced speech recognition abilities, our approach generates sentences that closely mirror the original text.
Figure 2: (a) Overview of the BrainECHO model framework. BrainECHO utilizes a three-stage training paradigm consisting of Mel spectrogram autoencoding, brain-audio latent space alignment, and Whisper finetuning. $C$, $T_{\varepsilon}$ denotes numbers of raw wave channels and timestamps, respectively. (b) Details of the Brain Encoder, which converts raw EEG/MEG signals into latent representations. $d$ represents the dimension of hidden states and TS Conv stands for Spatio-Temporal Convolution Networks. More details of Conformer are provided in Appendix A.
3 Method
3.1 Task Definition
@@
Given the raw EEG/MEG $E$, text content $T$, and corresponding audio stimuli $A$ during listening as mentioned in Section 4.1, the experimental data can be divided into a series of sentence-level EEG/MEG-text-speech pairs $\left\langle\varepsilon,t,a\right\rangle$. $\varepsilon\in\mathbb{R}^{C_{\varepsilon}\times T_{\varepsilon}}$, where $C_{\varepsilon}$ and $T_{\varepsilon}$ represent the channels and timestamps of brain signals, respectively. In general, $T_{\varepsilon}$ varies with the length of the sentence-level audio segment. Our goal is to decode the corresponding open-vocabulary tokens $t$ from the brain signal $\varepsilon$, with $a$ serving as auxiliary information.
3.2 Model Architecture
@@
Unlike the multi-task joint training employed in MAD (Yang et al., 2024c), BrainECHO adopts a three-stage training process. This method reduces resource consumption at each training step and facilitates the prediction of high-quality, high-resolution Mel spectrograms from brain signals. Specifically, we extend the spectrogram duration from 3 seconds, as used in  (Défossez et al., 2023; Yang et al., 2024c), to over 10 seconds, enabling sentence-level rather than segment-level brain-to-text translation, thereby preserving the semantics of the original sentences. The details of the model are shown in Figure 2. The following sections will detail each training stage.
3.2.1 Autoencoding of Audio Spectrogram
@@
Van den Oord et al. introduced the Vector Quantized-Variational AutoEncoder (VQ-VAE) (Van Den Oord et al., 2017) to learn discrete latent representations of audio, video, and other data types. Building on this approach, several studies (Li et al., 2023; Sadok et al., 2023; Yang et al., 2023) have explored representing Mel spectrograms using discrete tokens to capture phoneme-like information. Since Mel spectrograms effectively capture frequency and temporal patterns of audio, it is feasible to use them as an intermediate modality between brain signals and text (Metzger et al., 2023; Défossez et al., 2023). Due to the fact that the majority of existing strong audio autoencoders are pre-trained on audio waves rather than Mel, we chose to autoencode Mel spectrograms for obtaining a discrete representation space that is conducive to Mel reconstruction. Specifically, given a spectrogram $m\in\mathbb{R}^{T_{m}\times F_{m}}$, the audio encoder $Enc$ first converts it into a feature map $z_{m}=Enc(m)\in\mathbb{R}^{t_{m}\times f_{m}\times D}$ , where $T_{m}$, $F_{m}$ and $D$ denote the number of time frames, frequency bins and latent channels, respectively. The spectrogram is generated by the Whisper Processor, enabling text decoding from the reconstructed spectrogram using Whisper’s encoder-decoder architecture. Then, $z_{m}$ is processed by a vector quantizer $Q$. Specifically, each latent embedding $z_{m}^{ij}\in\mathbb{R}^{D}$ ($1\leq i\leq t_{m},1\leq j\leq f_{m}$) is replaced by the nearset vector $z_{q}^{ij}$ from a codebook $\mathbb{C}\in\mathbb{R}^{N\times D}$, which consists of $N$ learnable D-dimensional vectors. Formally, this process is expressed as follows:
@@
$\begin{split}Q(z_{m}^{ij})&=z_{q}^{ij}=c_{k},\\ \text{ where }k&=\mathop{\arg\min}\limits_{k\in\{1,2,...,N\}}\left\|z_{m}^{ij}-c_{k}\right\|_{2}.\end{split}$ (1)
@@
The reconstructed spectrogram is then obtained by the audio decoder $Dec$ as: $\hat{m}=Dec(z_{q})$. The encoder and decoder are both composed of ResUNet blocks (Kong et al., 2021). The training objective at this stage is defined as follows:
@@
3.2.2 Brain-Audio Latent Space Alignment
In the second stage, we freeze all the modules pre-trained in the previous stage and train a brain encoder to convert raw EEG/MEG signals $\varepsilon$ into latent features $z_{\varepsilon}$. The brain encoder utilizes a Conformer-based architecture (Song et al., 2022), which begins with Spatio-Temporal Convolutional Networks to process the input signals into a one-dimensional embedding sequence. The spatial convolutional layer reduces the number of input signal channels to one, while the temporal convolutional layers downsample the time dimension. This sequence is then added to learnable position embeddings and fed into a stack of Transformer encoder blocks. Linear layers and 2D convolutional networks subsequently transform the EEG/MEG features into representations matching the shape of $z_{m}$. Similarly, $z_{\varepsilon}$ is input into the frozen quantizer $Q$ and audio decoder $Dec$ to predict the corresponding Mel spectrogram $m$. Additionally, we align the representations of the Mel spectrogram and raw signals in the latent space. Notably, we employ a unified codebook to leverage pre-warmed discrete acoustic tokens for representing brain activity. Formally, the loss for stage 2 is as follows: